OpenStudy (lizz123):

@StudyGurl14 Identify the correct slope and y-intercept of the equation 5x - 3y = -6 A.slope = 5 over 3; y-intercept = (0, 2) B.slope = -5; y-intercept = (0, -6) C.slope = negative 3 over 5; y-intercept =(0,2) D.slope = 5 over 3; y-intercept = (0, -6)

OpenStudy (studygurl14):

Rewrite into slope-intercept form. 5x - 3y = -6 -3y = -5x - 6 y = (5/3)x + 2 In the form y = mx+ b m = slope b = y-intercpet

OpenStudy (anonymous):

First rewrite the equation in the form y=mx+c then the slope and y-intercept are m anc c respectively.@lizz123 The given equation can be written as 3y=5x+6 hence y=(5/3)x+2 Therefore slope=5/3 and y-intercept=(0,2) Option A is correct